OpenStudy (anonymous):

2x^2+2x=20 use thee discriminant to find the number of unique real solutions

OpenStudy (zzr0ck3r):

so 2x^2 + 2x -20=0 x^2 + 2 -10 = 0 -2 +- sqrt(2^2-4(1)(-10)))/(2a) make sence so far?

OpenStudy (anonymous):

yeah

OpenStudy (zzr0ck3r):

so the disceiminant is 2^2 - 4(1)(-10)

OpenStudy (zzr0ck3r):

discriminant > 0 2 real solutions discriminant = 0 1 real solution discriminant < 0 0 real soutions

OpenStudy (zzr0ck3r):

got it?

OpenStudy (anonymous):

yeah. Thank you

OpenStudy (anonymous):

By comparing the given equation with the standard: \[ax^2 + bx + c= 0\] You have a = 2, b =2 and c = -20 So formula for D is: \[D = b^2 = 4ac\] Plug in the values and find D..